If 3.6 grams of C react, how many moles of CS₂ will be produced?
1
Expert's answer
2020-05-01T14:42:42-0400

Let's write a balanced reaction equation:

C + 2S "\rightarrow" CS2

As you can see, when 1 mol of C is consumed, 1 mol of CS2 is produced. Therefore, the number of the moles of carbon and that of carbon sulfide are equal:

"n(C) = n(CS_2)"

The molar mass of carbon is 12.01 g/mol. Let's calculate the number of the moles of carbon:

"n(C) = \frac{m}{M} = \frac{3.6}{12.01} = 0.30" mol.

Therefore, 0.30 mol of CS2 is formed.

Answer: If 3.6 grams of C react, 0.30 mol of CS₂ will be produced.